OpenStudy (mattfort3212):

Which of the following expressions are polynomials? Check all that apply. A. x3 + 2x^2 B. 5/x-1 C. 6x^2+4sqr x D. -9x E. 2x^-3 - 5x^-1 + 3 F. 1/4x^5+1/3x^4-1/2x^3

sam (.sam.):

Polynomials, -poly "many", -nomial "terms". So, "Many terms" Examples of Polynomials: \[x+1,~~~~~3x^2+3x+2,~~~~~~\frac{2}{x}+1\] Examples of non-polynomials: \[x,~~x^2,~~3,~~\frac{2}{x-1}\]

OpenStudy (mattfort3212):

so it would be a b c e f

OpenStudy (mattfort3212):

no its a c e

sam (.sam.):

Nope, B and D are not polynomials

sam (.sam.):

What about F?

OpenStudy (mattfort3212):

i know d isnt so it acfe?

sam (.sam.):

Yup

OpenStudy (mattfort3212):

c is not a polynomial

sam (.sam.):

C is a polynomial

OpenStudy (mattfort3212):

it said a e f are and c was wrong.

sam (.sam.):

Look at it, it contains 2 terms \[6x^2+4\sqrt x\]

OpenStudy (mattfort3212):

yes but the quiz said no lol

sam (.sam.):

The quiz is wrong